OpenStudy (anonymous):

Simplify. (–24ab) ÷ (–3) A. -8ab B. -1/8ab C. 1/8ab D. 8ab

OpenStudy (anonymous):

Just do -24/-3= ? Answer this :)

OpenStudy (anonymous):

its D

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yea ! It s D . I wantes we dont use the answer :)

OpenStudy (anonymous):

can u help me with one more

OpenStudy (anonymous):

Sure :)

OpenStudy (austinl):

\((–24ab) ÷ (–3)=\dfrac{-24\times a\times b}{-3}\) \(\dfrac{-24}{-3}=8\) \(\large{\color{blue}{8ab}}\)